A steady current of 0.80 ampere was passed through an aqueous solution of tetraoxosulphate (6) acid for 7 hours. Calculate the volume of hydrogen gas that was liberated at standard temperature and pressure

Grade:10

1 Answers

Forum 1
one year ago
Quantity of electricity =Q=Ampere××Second
∴Q=0.8×420×60= 20160 Coulomb
Therefore, 11.2 dm3 ofH2 liberated by 96500 coulomb of electricity,
Thereforex dm3 ofH2liberated by 20160 coulomb of electricity .
∴ x=11.2 × 20160/96500= 2.34